Mal87:

If an object has a mass of 20 grams and a volume of 40 cm3, what is its density in g/cm3? A. 0.5 B. 2 C. 60 D. 800

dude:

Density=\(\large\frac{mass}{volume}\) Mass = 20g Volume = 40 \(cm^3\)

Mal87:

Thank you so much

dude:

Sure! Do you know what the density is now?

Mal87:

0.5

Mal87:

Answer A

mikewwe13:

20/40=0.5 g/cm^3 becuase, mass/volume=density.